OpenStudy (anonymous):

equation of the parabola wth axs parallel to y and passng thru (-1,-3) (1,-2),(2,1)

OpenStudy (anonymous):

general form for this parabola is y = ax^2 + bx + c plugging in the 3 points: -3 = a - b + c -2 = a + b + c 1 = 4a + 2b + c you need to solve this system of equations